# AcousTools: A ‘Full-Stack’, Python-Based, Acoustic Holography Library

### v6

@@ Abstract
Figure 1: The acoustic holography full-stack (top) describes the workflow for acoustic holography. A developer begins on the left and progresses to the right, moving from the most abstract stages (defining modelling parameters and modelling propagation) to the most physical stages (examining physical metrics of a sound field and producing that sound field in reality). This framework allows the analysis of different acoustic holography software based on the stages they support – the ideal software would allow a developer to complete every stage using a single software package. AcousTools uniquely addresses all stages in the full stack as shown in this example (bottom), from (1) defining the board and scatterer for levitation setup (i.e., a flat array and rabbit reflector), (2) computing the sound propagation using the scattering profile, (3) solving for the transducers’ activation to focus sound waves above the rabbit’s head, and then (4) analysing the acoustic force on a small particle (shown as blue arrows) before (5) finally rendering this in the real-world hardware.
@@ 1 Introduction
Non-contact manipulation and stimulus open the door for numerous applications in a user-friendly and ergonomic way. Objects can be fabricated without ever touching the materials used, either by assembling smaller building blocks [LeviPrint] or by shaping raw materials [AcoustoFab]. Haptic stimulus can give greater immersion in virtual reality [HapticVR], leveraging human tactile mechanisms to produce higher quality stimulus [HapticsMPSTM, HapticIllusions] and render complex stimulus such as softness [HapticSoftness]. Levitation of small particles, in addition to fabrication, allows the creation of computer displays that are truly 3D [VolDisplaySingleParticle], by using props to represent objects [Articulev] or by exploiting persistence of vision (PoV) to render shapes in 3D [GSPAT, VolDisplaySingleParticle]. All these applications and more are unlocked by utilising focused mid-air ultrasound.
To create these applications, a software package or library that can model the behaviour of acoustic waves is advantageous [OpenMPD, Ultraino]. Common tasks include computing phases for transducer arrays in complex scenarios [LeviPrint], predicting how sound scatters from irregular objects [HighSpeedBEM], and analysing the force exerted on an object in a sound field [AcousticElements]. Pre-built software allows developers to focus on higher-level goals without needing to implement these complex processes from scratch.
@@
There are a variety of software packages that attempt to meet the stages described in the full-stack [OpenMPD, Ultraino, Levitate, Kwave, FOCUS]. However, all existing solutions fail to meet one or many of these categories as can be seen in Tab.1. As a result, developers often need to switch between different packages or are unable to use certain methodologies without implementing them independently.
@@
In this work, we step through each stage of the full-stack, showing AcousTools’ efficacy in that area and highlighting where other software falls short. Its implementation using PyTorch in Python leverages the language’s ease of access especially for beginners [pythonEase], as well as the parallelised and comprehensive capabilities of the PyTorch environment [PyTorch]. AcousTools is therefore the general purpose software solution that the acoustic holography community currently lacks. We anticipate that the community will adopt it as a standard code library, leading to a greater degree of interoperability and reproducibility among researchers.
@@ 2.1 Acoustic Holography
Acoustic holography, used to produce 3D sound fields[HoloForAcoustics, AcousticElements, GSPAT, HighSpeedBEM], can be used to focus high frequency sound waves. This allows for non-contactless manipulation [VolDisplaySingleParticle, LeviPrint] and stimulation [eigensolver, HapticsMPSTM] of materials at a distance. Unlike magnetic levitation, it is not limited to ferrous materials, and unlike optical holography using lasers, there is no dangerous heating of the samples [2DDroplets]. Thus, acoustic holography is an exciting and interactive method. Acoustic levitation is even used in the field of nanoparticle mid-air chemistry [AcousticChemistry], but only the simplest acoustic setups [TinyLev] are used in these cases due to the complexity of the control software commonly used in holographic work. Modern open-source hardware[Ultraino, OpenMPD] combined with reliable and flexible software has a potential to expand the domains where acoustic holography can be used.
Propagation Methods: Acoustic holography computes a set of transducer activations used to produce a desired 3D sound field [AcousticElements]. It predicts how these waves interfere and build up [VolDisplaySingleParticle, GSPAT, Diff-PAT] by modelling sound propagation. Sound waves are emitted from a source, often an array of transducers, and typically modelled using the Piston Model [GSPAT, LeviPrint, HapticsMPSTM, VolDisplaySingleParticle, AcoustoFab]. In this model, the propagation from a single transducer $t\in\boldsymbol{\tau}$, for a set of transducers $\boldsymbol{\tau}$, located at $\boldsymbol{z}_{t}$ to a point in space located at $\boldsymbol{z}_{n}$ can be modelled as [GSPAT]
@@
In a free-field scenario without scattering or occlusion, this model is sufficient. However, considering scattering would provide benefits to many applications. When there are objects in a levitation domain, for example when an object is fabricated during levitation, modelling sound scattering can increase manipulation precision and reduce potential levitation failure [AcoustoFab]. This can be done by using a three step model, known as the boundary element method (BEM). Instead of the free-field propagator $\boldsymbol{F}_{\boldsymbol{\nu}\boldsymbol{\tau}}$, this model uses an extended propagator $\boldsymbol{E}_{\boldsymbol{\nu}\boldsymbol{\tau}}$, which is formed as [HighSpeedBEM]
@@
Where $\boldsymbol{F}_{\boldsymbol{\nu}\boldsymbol{\tau}}$ is the same as in (1), $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ models the propagation of the sound from the transducers to a mesh, $\boldsymbol{\mu}$, representing a scattering object, and $\boldsymbol{G}_{\boldsymbol{\nu}\boldsymbol{\mu}}$ models how the waves travel from the mesh surface to the points. Thus, the overall propagator $\boldsymbol{E}_{\boldsymbol{\nu}\boldsymbol{\tau}}$ considers both the direct ($\boldsymbol{F}_{\boldsymbol{\nu}\boldsymbol{\tau}}$) and the scattered ($\boldsymbol{G}_{\boldsymbol{\nu}\boldsymbol{\mu}}\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$) components. However, while using BEM improves many applications[HighSpeedBEM, ToBEMorNot], the increased computational complexity, especially in computing $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ holds back its utility [ToBEMorNot]. As $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ depends only on the geometry of the scattering objects, it can be precomputed and cached to reduce this cost in some cases [HighSpeedBEM]. However, if the object’s geometry changes, $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ must be recomputed, and so dynamic scenarios will suffer from the increase in computation speeds[ToBEMorNot].
@@
Phase Retrieval Algorithms: For a given propagator, the transducer activations $\boldsymbol{x}$, which we refer to as an acoustic hologram, need to be determined. A generalised objective function can be defined as [TGS]
@@
Here, $\mathcal{L}$ is an objective function of the complex field produced by the propagation of the hologram x, for example maximising the pressure amplitude at the targets[IB, GSPAT, TGS]. The constraint $\mathcal{C}$ quantifies hardware requirements for the hologram x, often using indicator functions that are zero if these requirements are met and infinite otherwise. Common constraints include forcing transducer amplitudes to be maximum [IB], $|\textbf{{x}}|=1$, or allowing amplitudes to vary within physically feasible limits [GSPAT], $|\textbf{{x}}|\leq 1$. Methods to fulfil (4) take two main forms, gradient-descent as well as projective methods.
@@
In order to then perform this optimisation, computing the gradients can be done either explicitly or implicitly. Explicit computations require more labour from the developer [AcousticElements], where they must directly know and implement how to find the gradients. Conversely, implicit gradients are more convenient and versatile [Diff-PAT, Diff-PAT-Expanded] as the exact form of the gradients are avoided by utilising automatic differentiation. This accumulates the gradients through computation of an objective and then uses the chain rule to backpropagate through this computation tree to compute the gradients of a wide variety of functions [Diff-PAT, Diff-PAT-Expanded]. This gives flexibility in algorithm design as the gradient to minimise an objective need not to be defined directly. However, to use automatic differentiation, the chain of operations needs to be maintained, requiring specific implementations such as that given in the Python packages PyTorch[PyTorch] or Jax [jax].
On the other hand, alternating projections are based on – normally iterative – algorithms, in which forward and backward propagations between the ultrasound source and the target points are combined with projections to feasible values. The most common algorithms are based on the Gerchberg-Saxton (GS) algorithm[GS], first described in acoustics as the iterative backpropagation algorithm (IB), which can be written as
@@
In this case, $\mathcal{P}_{\Pi}$ is introduced as a second indicator constraint that sets the point amplitudes to some target point amplitude y, defined as $\mathcal{P}_{\Pi}(\textbf{{p}}|\boldsymbol{y})=(\textbf{{p}}/|\textbf{{p}}|)\cdot\boldsymbol{y}$ given the point activation p and element wise multiplication $(\cdot)$. $\boldsymbol{A}$ propagates the hologram to the target points pressure, this is then constrained by $\mathcal{P}_{\Pi}$. This constrained point is back-propagated using the conjugate transpose of the propagator, $\boldsymbol{A}^{*}$, which is itself constrained by $\mathcal{P}_{\chi}$. The IB algorithm is known to act as projected gradient descent for an objective function that maximizes the sum of amplitudes at target points[TGS].
@@
This results in very fast computation rates, but failing to finely equalize pressure at target points. Alternatively, GS-PAT removes the target points amplitude constraints, combining forward and backward propagation for high computational speed [GSPAT], while Weighted Gerchberg-Saxton (WGS) weights the pressure at each target point during each iteration to achieve low variance in pressure across target points[TGS]. Nevertheless, beyond interpretability, alternating projection algorithms also lack gradient descent’s adaptability to adapt to custom objectives that may be required by users for different applications.
Gor’kov potential, Force, and Stiffness: When the field has been modelled for a given set of transducer activations, various metrics can be computed that quantify the field. The simplest measure of an acoustic field is the pressure at a point in space. When the solver has correctly focused waves at a point, there should be a local maxima of pressure at that point and the amplitude of the sound field at position can be described by $|\boldsymbol{A}\boldsymbol{x}|$. Similarly, the phase can be found by taking the complex argument of $\boldsymbol{A}\boldsymbol{x}$.
The Gor’kov potential is a potential field based on both the pressure and gradient of pressure surrounding a small particle as [AcousticElements]
@@
Here, $c_{0}$ and $c_{p}$ are the speed of sound in the medium and the particle, $\rho_{0}$ and $\rho_{p}$ are the density of the medium and particle, $\omega$ is the angular frequency, and $V$ is the volume of the particle. $p$ is the pressure at a point and ${\partial p}/{\partial a}$ is the gradient of pressure in the $a$-axis where $a\in(x,y,z)$ [AcousticElements]. The Gor’kov potential is the field such that the negative gradient of the Gor’kov potential is equal to the force acting on a small particle in the sound field.
@@
As a result, the point at which a particle will levitate corresponds to where the gradient of Gor’kov potential is zero (so there is no net force at that point). In reality, the levitation location will be very slightly below this point due to the force of gravity but this difference can be largely neglected for very small particles [ReviewOfProgress]. The points with $\boldsymbol{\nabla}U=0$ includes both minima and maxima of Gor’kov potential and both are technically valid levitation locations (with net force equal to zero). However, at a maxima of Gor’kov potential, the forces diverge, and therefore the trap becomes unstable. To quantify a stable trap, we can use the stiffness at a point in space, which can be represented as the sum of the unmixed, second gradients (the Laplacian) of the Gor’kov potential as [AcousticElements]
@@ 2.2 Existing Software
AcousTools OpenMPD[OpenMPD] Ultraino[Ultraino] Levitate[Levitate]
@@
Propagators Piston Model [GSPAT] Y Y Y Y
Propagators BEM [HighSpeedBEM] Y N N N
Solvers Naive [VolDisplaySingleParticle] Y Y Y N
Solvers IB[IB] Y Y N N
Solvers GS-PAT[GSPAT] Y Y N N
Solvers Gradient Descent[AcousticElements] Y N N Y
Solvers Implicit-Gradient [Diff-PAT] Y N N N
@@
Analysis Gor’kov Potential [AcousticElements] Y Y Y Y
Analysis Force [AcousticElements] Y Y Y Y
Analysis Stiffness [AcousticElements] Y Y N Y

### v7

@@ Abstract
Figure 1: The acoustic holography full-stack (top) describes the workflow for acoustic holography. A developer begins on the left and progresses to the right, moving from the most abstract stages (defining modelling parameters and modelling propagation) to the most physical stages (examining physical metrics of a sound field and producing that sound field in reality). This framework allows the analysis of different acoustic holography software based on the stages they support – the ideal software would allow a developer to complete every stage using a single software package. AcousTools uniquely addresses all stages in the full stack as shown in this example (bottom), from (1) defining the board and scatterer for levitation setup. Here a flat array above the working volume facing down towards a flat surface made of acrylic with an acrylic rabbit on top, together acting as a reflector. The array and reflector are separated by 12cm of air, (2) computing how the sound propagates with the scattering profile by using the boundary element method, (3) solving for the transducers’ activation to focus sound waves above the rabbit’s head while accounting for scattering, and then (4) analysing the acoustic force on a small particle (shown as blue arrows) before (5) finally rendering this in the real-world hardware.
@@ 1 Introduction
Non-contact manipulation and stimulus opens the door for numerous applications in a user-friendly and ergonomic way. Objects can be fabricated without ever touching the materials used, either by assembling smaller building blocks [7] or by shaping raw materials [5]. Haptic stimulus can give greater immersion in virtual reality [14], leveraging human tactile mechanisms to produce higher quality stimulus [45, 44] and render complex stimulus such as softness [33]. Levitation of small particles, in addition to fabrication, allows the creation of computer displays that are truly 3D [19], by using props to represent objects [8] or by exploiting persistence of vision (PoV) to render shapes in 3D [41, 19]. All these applications and more are unlocked by utilising focused mid-air ultrasound.
To create these applications, a software package or library that can model the behaviour of acoustic waves is advantageous [36, 30]. Common tasks include computing phases for transducer arrays in complex scenarios [7], predicting how sound scatters from irregular objects [18], and analysing the force exerted on an object in a sound field [32]. Pre-built software allows developers to focus on higher-level goals without needing to implement these complex processes from scratch.
@@
There are a variety of software packages that attempt to meet the stages described in the full-stack [36, 30, 1, 48, 9]. However, all existing solutions fail to meet one or many of these categories as can be seen in Tab.1. As a result, developers often need to switch between different packages or are unable to use certain methodologies without implementing them independently.
@@
In this work, we step through each stage of the full-stack, showing AcousTools’ efficacy in that area and highlighting where other software falls short. Its implementation using PyTorch in Python leverages the language’s ease of access especially for beginners [47], as well as the parallelised and comprehensive capabilities of the PyTorch environment [40]. AcousTools is therefore the general purpose software solution that the acoustic holography community currently lacks. We anticipate that the community will adopt it as a standard code library, leading to a greater degree of interoperability and reproducibility among researchers.
@@ 2.1 Acoustic Holography
Acoustic holography, used to produce 3D sound fields [34, 32, 41, 18], can be used to focus high frequency sound waves. This allows for non-contactless manipulation [19, 7] and stimulation [28, 45] of materials at a distance. Unlike magnetic levitation, it is not limited to ferrous materials, and unlike optical holography using lasers, there is no dangerous heating of the samples [10]. Thus, acoustic holography is an exciting and interactive method. Acoustic levitation is even used in the field of nanoparticle mid-air chemistry [42], but only the simplest acoustic setups [29] are used in these cases due to the complexity of the control software commonly used in holographic work. Modern open-source hardware [30, 36] combined with reliable and flexible software has a potential to expand the domains where acoustic holography can be used.
Propagation Methods: Acoustic holography computes a set of transducer activations used to produce a desired 3D sound field [32]. It predicts how these waves interfere and build up [19, 41, 11] by modelling sound propagation. Sound waves are emitted from a source, often an array of transducers, and typically modelled using the Piston Model [41, 7, 45, 19, 5]. In this model, the propagation from a single transducer $t\in\boldsymbol{\tau}$, for a set of transducers $\boldsymbol{\tau}$, located at $\boldsymbol{z}_{t}$ to a point in space located at $\boldsymbol{z}_{n}$ can be modelled as [41]
@@
In a free-field scenario without scattering or occlusion, this model is sufficient. However, considering scattering would provide benefits to many applications. When there are objects in a levitation domain, for example when an object is fabricated during levitation, a hand is present when rendering haptics or a barrier blocks a levitation path, modelling sound scattering can increase manipulation precision and reduce potential levitation failure [5]. This scattering media can be any sound-hard material, such as plastics or the surface of water and as many common materials have an acoustic impedance much higher than that of air, most materials act as scatterers when using air as a medium [18].
Predicting the scattering behaviour can be conducted by using a three step model, known as the boundary element method (BEM). Instead of the free-field propagator $\boldsymbol{F}_{\boldsymbol{\nu}\boldsymbol{\tau}}$, this model uses an extended propagator $\boldsymbol{E}_{\boldsymbol{\nu}\boldsymbol{\tau}}$, which is formed as [18]
@@
Where $\boldsymbol{F}_{\boldsymbol{\nu}\boldsymbol{\tau}}$ is the same as in (1), $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ models the propagation of the sound from the transducers to a mesh, $\boldsymbol{\mu}$, representing a scattering object, and $\boldsymbol{G}_{\boldsymbol{\nu}\boldsymbol{\mu}}$ models how the waves travel from the mesh surface to the points. Thus, the overall propagator $\boldsymbol{E}_{\boldsymbol{\nu}\boldsymbol{\tau}}$ considers both the direct ($\boldsymbol{F}_{\boldsymbol{\nu}\boldsymbol{\tau}}$) and the scattered ($\boldsymbol{G}_{\boldsymbol{\nu}\boldsymbol{\mu}}\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$) components. However, while using BEM improves many applications [18, 37], the increased computational complexity, especially in computing $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ holds back its utility [37]. Both $\boldsymbol{G}_{\boldsymbol{\nu}\boldsymbol{\mu}}$ and $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ have the number of mesh elements as a dimension, and as such a mesh with many elements per wavelength (in order to improve model accuracy) will significantly increase the size of these matrices. Although, as $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ depends only on the geometry of the scattering objects, it can be precomputed and cached to reduce this cost in some cases [18]. However, if the object’s geometry changes, $\boldsymbol{H}_{\boldsymbol{\mu}\boldsymbol{\tau}}$ must be recomputed, and so dynamic scenarios will suffer from the increase in computation speeds [37].
@@
Phase Retrieval Algorithms: For a given propagator, the transducer activations $\boldsymbol{x}$, which we refer to as an acoustic hologram, need to be determined. A generalised objective function can be defined as [6]
@@
Here, $\mathcal{L}$ is an objective function of the complex field produced by the propagation of the hologram x, for example maximising the pressure amplitude at the targets [31, 41, 6]. The constraint $\mathcal{C}$ quantifies hardware requirements for the hologram $\boldsymbol{x}$, often using indicator functions that are zero if these requirements are met and infinite otherwise. Common constraints include forcing transducer amplitudes to be maximum [31], $|\textbf{{x}}|=1$, or allowing amplitudes to vary within physically feasible limits [41], $|\textbf{{x}}|\leq 1$. Methods to fulfil (4) take two main forms, gradient-descent as well as projective methods.
@@
In order to then perform this optimisation, computing the gradients can be done either explicitly or implicitly. Explicit computations require more labour from the developer [32], where they must directly know and implement how to find the gradients. Conversely, implicit gradients are more convenient and versatile [11, 12] as the exact form of the gradients are avoided by utilising automatic differentiation. This accumulates the gradients through computation of an objective and then uses the chain rule to backpropagate through this computation tree to compute the gradients of a wide variety of functions [11, 12]. This gives flexibility in algorithm design as the gradient to minimise an objective need not to be defined directly. However, to use automatic differentiation, the chain of operations needs to be maintained, requiring specific implementations such as that given in the Python packages PyTorch [40] or Jax [21].
On the other hand, alternating projections are based on – normally iterative – algorithms, in which forward and backward propagations between the ultrasound source and the target points are combined with projections to feasible values. The most common algorithms are based on the Gerchberg-Saxton (GS) algorithm [15], first described in acoustics as the iterative backpropagation algorithm (IB), which can be written as
@@
In this case, $\mathcal{P}_{\Pi}$ is introduced as a second indicator constraint that sets the point amplitudes to some target point amplitude y, defined as $\mathcal{P}_{\Pi}(\textbf{{p}}|\boldsymbol{y})=(\textbf{{p}}/|\textbf{{p}}|)\cdot\boldsymbol{y}$ given the point activation p and element wise multiplication $(\cdot)$. $\boldsymbol{A}$ propagates the hologram to the target points pressure, this is then constrained by $\mathcal{P}_{\Pi}$. This constrained point is back-propagated using the conjugate transpose of the propagator, $\boldsymbol{A}^{*}$, which is itself constrained by $\mathcal{P}_{\chi}$. The IB algorithm is known to act as projected gradient descent for an objective function that maximizes the sum of amplitudes at target points [6].
@@
This results in very fast computation rates, but failing to finely equalize pressure at target points. Alternatively, GS-PAT removes the target points amplitude constraints, combining forward and backward propagation for high computational speed [41], while Weighted Gerchberg-Saxton (WGS) weights the pressure at each target point during each iteration to achieve low variance in pressure across target points [6]. Nevertheless, beyond interpretability, alternating projection algorithms also lack gradient descent’s adaptability to adapt to custom objectives that may be required by users for different applications.
Alternatively, recent developments in machine learning (ML) been applied to phase-retrieval tasks for acoustic holography. For example, AcousNet which uses a convolutional neural network (CNN) in order to generate image holograms much faster than some iterative algorithms such as IB could [51]. In a similar way, HU-Net [27] and EIR-Net [49] both predict the transducer signals based off a target plane with EIR-Net directly improving on the performance of AcousNet [49]. ML has also been applied in the context of MHz-imaging ultrasound with the Ultrasound Deep Learning Model achieving close and well defined focal points [43].
Gor’kov potential, Force, and Stiffness: When the field has been modelled for a given set of transducer activations, various metrics can be computed that quantify the field. The simplest measure of an acoustic field is the pressure at a point in space. When the solver has focused waves at a point, there should be a local maxima of pressure at that point and the amplitude of the sound field at position can be described by $|\boldsymbol{A}\boldsymbol{x}|$. Similarly, the phase can be found by taking the complex argument of $\boldsymbol{A}\boldsymbol{x}$.
The Gor’kov potential is a potential field based on both the pressure and gradient of pressure surrounding a small particle as [32]
@@
Here, $c_{0}$ and $c_{p}$ are the speed of sound in the medium and the particle, $\rho_{0}$ and $\rho_{p}$ are the density of the medium and particle, $\omega$ is the angular frequency, and $V$ is the volume of the particle. $p$ is the pressure at a point and ${\partial p}/{\partial a}$ is the gradient of pressure in the $a$-axis where $a\in(x,y,z)$ [32]. The Gor’kov potential is the field such that the negative gradient of the Gor’kov potential is equal to the force acting on a small particle in the sound field.
@@
As a result, the point at which a particle will levitate corresponds to where the gradient of Gor’kov potential is zero (so there is no net force at that point). In reality, the levitation location will be very slightly below this point due to the force of gravity but this difference can be largely neglected for very small particles [2]. The points with $\boldsymbol{\nabla}U=0$ includes both minima and maxima of Gor’kov potential and both are technically valid levitation locations (with net force equal to zero). However, at a maxima of Gor’kov potential, the forces diverge, and therefore the trap becomes unstable. To quantify a stable trap, we can use the stiffness at a point in space, which can be represented as the sum of the unmixed, second gradients (the Laplacian) of the Gor’kov potential as [32]
@@ 2.2 Existing Software
AcousTools OpenMPD [36] Ultraino [30] Levitate [1]
@@
Propagators Piston Model [41] Y Y Y Y
Propagators BEM [18] Y N N N
Solvers Naive [19] Y Y Y N
Solvers IB [31] Y Y N N
Solvers GS-PAT [41] Y Y N N
Solvers Gradient Descent [32] Y N N Y
Solvers Implicit-Gradient [11] Y N N N
@@
Analysis Gor’kov Potential [32] Y Y Y Y
Analysis Force [32] Y Y Y Y
Analysis Stiffness [32] Y Y N Y